5x + 4y = 40
x + 2y = 11

Answer by mananth(16946)

x+4y=40 .............1
x+2y=11 .............2

multiply (1)by 1
Multiply (2) by -2
x+4y=40
-2x-4y=-22

Add the two equations
x=-18

plug value ofxin (1)

x+4y=40
-18+4y=40
4y=40+18
4y=58
y=14.50